# Nwāchā Munā: A Devanagari Speech Corpus and Proximal Transfer Benchmark for Nepal Bhasha ASR

### 08 Mar 2026 v1

@@ Abstract
Nepal Bhasha (Newari), an endangered language of the Kathmandu Valley, remains digitally marginalized due to the severe scarcity of annotated speech resources. In this work, we introduce Nwāchā Munā (/nwa:tSa: muna:/), a newly curated 5.39-hour manually transcribed Devanagari speech corpus for Nepal Bhasha, and establish the first benchmark using script-preserving acoustic modeling. We investigate whether proximal cross-lingual transfer from a geographically and linguistically adjacent language (Nepali) can rival large-scale multilingual pretraining in an ultra-low-resource Automatic Speech Recognition (ASR) setting. Fine-tuning a Nepali Conformer model reduces the Character Error Rate (CER) from a 52.54% zero-shot baseline to 17.59% with data augmentation, effectively matching the performance of the multilingual Whisper-Small model despite utilizing significantly fewer parameters. Our findings demonstrate that proximal transfer within South Asian language clusters serves as a computationally efficient alternative to massive multilingual models. We openly release the dataset and benchmarks to digitally enable the Newari community and foster further research in Nepal Bhasha.
Keywords:  Automatic Speech Recognition, Nepal Bhasha, Speech Corpus, Low-Resource Languages
@@
Rupak Tiwari, Arju Shrestha,
Rupak Raj Ghimire, Bal Krishna Bal
@@ 1.  Introduction
In the current space of AI, high-resource languages dominate in terms of resource availability while indigenous languages remain digitally marginalized. In a country like Nepal, with 124 official languages, this divide is especially vivid, creating a need for technologies like ASR to bridge the gap. Nepal Bhasha (Newari) serves as a prime example of this imbalance, despite being actively spoken by over 860,000 people in Nepal. Even with a six-century old history and a designation as one of the official working languages of Bagmati Province Post Report (2024), Nepal Bhasha is currently classified by UNESCO as a ``definitely endangered'' language Moseley (2010). Building a system such as ASR empowers these communities to maintain their linguistic heritage while fully participating in the modern technological landscape.
@@
This work addresses the following research question:
Can proximal transfer from a geographically and linguistically related South Asian language provide performance comparable to massive multilingual pretraining in an ultra-low-resource ASR setting?
By answering this, we aim to provide a scalable blueprint for other underrepresented languages in the region.
@@ 2.  Related works
One major obstacle in building ASR systems for underrepresented languages is the lack of sufficient speech data. Older methods like Hidden Markov Models required less data but struggled to capture intricate acoustic patterns. With time, new architectures like Transformers Vaswani et al. (2017) and Conformer Gulati et al. (2020) have outperformed older statistical methods, but they require massive datasets. To address this issue, researchers have increasingly turned to transfer learning. For instance, Cheng et al. (2024) demonstrated clear ASR performance gains for two at-risk Austronesian languages, Amis and Seediq, despite having minimal available speech samples. They achieved this by utilizing a novel data-selection scheme to automatically extract phonetically similar utterances from a broad multilingual corpus to augment their training data Likewise, significant progress has been made specifically within the South Asian region. The AI4Bharat initiative Javed et al. (2022) advanced regional ASR by curating over 17,000 hours of raw speech data across 40 Indian languages for continued pre-training.
@@
In the context of Nepali ASR, early efforts relied on Hidden Markov Models (HMM) Ssarma et al. (2017) for isolated word recognition, which struggled with continuous speech. The transition to deep learning began with Regmi et al. (2019), who implemented Recurrent Neural Networks (RNN) with Connectionist Temporal Classification (CTC) loss. This was further refined by Bhatta et al. (2020), who introduced a hybrid CNN-GRU-CTC model. To address training instability in deeper networks, Dhakal et al. (2022) proposed a CNN-BiLSTM-ResNet architecture, achieving a CER of 17.06%. Concurrently, research began shifting towards data efficiency and cross-lingual techniques; Bansal et al. (2020) explored an acoustic-phonetic approach, demonstrating that leveraging monolingual and cross-lingual information could improve performance in low-resource settings. This era also saw the first steps towards attention mechanisms, with Joshi and Shrestha (2023) employing Self-Attention Networks (SAN) to capture long-range dependencies better than recurrent models.
@@
Current state-of-the-art systems have fully embraced E2E Transformer-based architectures. Poudel et al. (2025) introduced NepConformer, a Conformer-based Nepali ASR model achieving a CER of 6.01%. Recent focus has also turned to data selection, tokenization and resource efficient strategies. Ghimire et al. (2023a) proposed an Active Learning framework to select the most informative samples for speech annotation. Ghimire et al. (2023b) demonstrated that using linguistically motivated sub-word units (syllables) alongside active sampling significantly outperforms standard methods. Ghimire et al. (2025) showed that using adapter-based Parameter-Efficient-Fine-Tuning (PEFT), Word Error Rate (WER) improves by 19% compared to fine-tuning Whisper-v2. Most recently, Ghimire et al. (2023b) addressed grammatical consistency by proposing a Rule-Based Character Constituency Loss (RBCCL), which reduced Word Error Rates (WER) from 47.1% to 23.41% by penalizing grammatically impossible character sequences during training.
@@
While Nepali has seen consistent growth in the field of ASR, research on Nepal Bhasha remains sparse. A notable exception is the work by Meelen et al. (2024), who presented the first Newari ASR. where they achieved a CER of 12%. Their evaluation is conducted in Romanized transliteration, whereas our work preserves Devanagari orthography. Due to differences in script representation, normalization, and evaluation protocol, direct numerical comparison must be interpreted cautiously. Their study highlighted that while transfer learning is effective, the lack of standardized orthography and limited data results in higher WER compared to Nepali. They emphasized the need for "orthography standardization" and data augmentation techniques like SpecAugment Park et al. (2019) to make ASR viable for these endangered languages.
@@
Taken together, these findings highlight a persistent gap: prior work lacks ASR systems trained on the Devanagari script using carefully curated low-resource datasets. This research clears these gaps by developing Nwāchā Munā. We utilise Devnagari script for transcribing speech of Nepal Bhasha. We also use data augmentation techniques to train robust ASR models using various cross-lingual transfer strategies and report CER for each technique.
@@ 3.1.  Data Collection Technique
The development of the Nepal Bhasha speech recognition system required a carefully designed data collection process to ensure authenticity and reliability. Both textual and acoustic resources were gathered considering language and speaker diversity. The objective was to construct a dataset that reflects natural language usage while maintaining consistency and structural integrity. The following subsections describe the procedures adopted for textual and audio data acquisition.
@@ 3.1.1.  Textual Data Acquisition
To establish a baseline for the Nepal Bhasha ASR system, a comprehensive textual corpus was compiled from a different sources of digital and print media. The data collection strategy included the use of both formal and everyday language to capture the full range of vocabulary. Digital data were extracted from Nepal Bhasha Wikipedia 1 and the OSCAR (Open Super-large Crawler-Aisled) Ortiz Suárez et al. (2020) dataset while print-based resources included regional newspapers, literary manuscripts from the Bloom Library 2, and primary school textbooks. In addition to the formal texts obtained, a targeted set of daily conversational sentences was curated manually to incorporate texts that were previously missing. The raw text was rigorously pre-processed where non-target language tokens, numerical digits, and idiosyncratic symbols were manually removed to ensure high fidelity and relevance. The overall text corpus statistics and sentence length distribution are presented in Table 1 and 2:
@@
Total word-level tokens 27,644
Unique word in corpus 8,599
@@ 3.1.2.  Audio Data Acquisition
Primary acoustic data was gathered through a dual-modal strategy involving original field recordings and the transcription of existing web-based audio, ensuring the model's exposure to varied acoustic environments. To maintain technical consistency across the dataset, all audio was standardized to a 16 KHz sampling rate and stored in a mono-channel WAV format, providing the necessary fidelity for deep learning-based feature extraction. For the primary recording phase, 18 native speakers from the community in Banepa, Dhulikhel, Panauti and Patan were recruited, with selection criteria like native fluency and linguistic knowledge. Since most of the indigenous Newar population concentrate around the Kathmandu Valley (National Statistics Office, 2023), data collection was primarily focused on these specific areas. To enhance the model's robustness against speaker-dependent variability, the data was demographically stratified by age and gender. Recording was performed in an area with minimal acoustic disturbance. This field collection of 4 hours and 21 minutes of speech was further supplemented by approximately one hour of web-sourced audio (Meelen et al., 2024), which was originally in Romanized form and subsequently transliterated into Devanagari script by community members. The overall characteristics of the NewaASR speech corpus are summarized in Tables 3–5. Table 3 presents the corpus-level statistics, including the total number of utterances, total duration, and mean utterance length. Table 4 details the age distribution of the same participant set, while Table 5 provides the gender distribution of the 18 speakers contributing to the corpus.
@@ 3.2.1.  Acoustic Modeling Strategies
We employed augmentation strategies tailored to each architecture. For the Conformer model, we implemented static augmentation, physically expanding the training set by a factor of 5 via offline speed perturbation ($0.9\times,1.1\times$), volume randomization, and noise injection. Likewise, a dynamic data augmentation was implemented before model input, audio samples were subjected to random transformations including time stretching (rate $0.8\times$$1.25\times$), pitch shifting ($\pm 4$ semitones), and Gaussian noise injection (amplitude 0.0010.0010.0150.015), each applied with a probability of $p=0.5$. This strategy effectively regularized the model by ensuring it encountered a diverse range of acoustic variations, preventing memorization of the limited training set.
@@ 3.2.2.  Language Modeling and Decoding
In the decoder-only adaptation setup, the encoder parameters were frozen, and only the prediction network and joint network were updated during fine-tuning. By applying the phonetic similarities between Nepali and Newari, this strategy conserves the transferable acoustic representations obtained during pre-training while allowing the decoder to adapt specifically to Newari linguistic patterns.
Independent of this architectural adaptation, we also integrated an external n-gram language model trained using KenLM and applied shallow fusion during inference. This introduces explicit linguistic priors to capture word-level dependencies that the limited paired data might miss. During beam search decoding, the log-probabilities from the acoustic model and the external language model were linearly combined (Toshniwal et al., 2018), with the fusion weight tuned on the development set.
For the baseline system, beam search decoding approximates the following objective:
@@
$y^{*}=\arg\max_{y}\left\{\log p(y\mid x)+\lambda\log p_{LM}(y)+\beta|y|_{w}\right\}$ (2)
where $\lambda$ controls the language model weight, $\beta$ is the word insertion bonus, and $|y|_{w}$ denotes the number of words in the hypothesis $y$. In this work, we utilized a 5-gram KenLM with $\lambda=0.4$ and $\beta=1.5$, tuned on the development set using a beam width of 128.
Decoder-only fine-tuning improves the model’s language understanding by updating only the higher-level components while keeping the encoder fixed, which also reduces computational cost. This setup allows us to observe the effect of adapting the decoder alone in a low-resource setting. In a separate experiment, shallow fusion adds an external language model during decoding without changing the acoustic model itself. By evaluating these two methods independently, we can clearly measure the contributions of internal decoder adaptation and external statistical language modeling to CER, WER, and improve transcription consistency in low-resource settings.
@@ 3.2.3.  Semi-Supervised Learning
Due to the availability of limited data of Nepal Bhasha, we adopted a semi-supervised pseudo-labeling framework. An unlabeled corpus of 13.65 hours was collected from publicly available broadcast sources, including radio, podcasts, and television, and segmented into shorter utterances. To mitigate memory issues and alignment degradation in CTC-based training, utterances longer than 15 seconds were discarded. The remaining audio was transcribed using our best intermediate acoustic model to generate pseudo-labels.
To ensure quality, we applied rigorous filtering. Single-word transcripts were removed, as they were typically noise-induced fragments rather than valid words. We also computed average prediction confidence scores and applied a 70% threshold to reduce error reinforcement. After filtering, 9.33 hours of high-quality pseudo-labeled data remained.
However, incorporating this data with the original labeled training set degraded performance, increasing CER to 19.83% compared to the 17.59% baseline. The decline in CER highlights a critical challenge in ultra-low-resource settings: pseudo-labeling from acoustically diverse broadcast data may introduce domain shift that outweighs gains from increased data volume. This suggests that domain alignment may be more important than raw quantity in endangered-language ASR, as uncurated broadcast sources can introduce background noise and out-of-domain speeches which negatively affects the acoustic modeling despite confidence-based filtering.
@@ 4.1.  Experimental Setup
The fine-tuned Newari ASR model utilizes a Conformer-CTC architecture with a Byte Pair Encoding (BPE) tokenizer of vocabulary size 128. The 18-layer Conformer encoder has a d-model dimension of 256, 4 attention heads, and a convolution kernel size of 31. Regularization includes a dropout rate of 0.1 across the encoder and attention modules, alongside SpecAugment with 2 frequency masks (width 27) and 2 time masks (width 70). Optimization is performed using AdamW with a learning rate of $1\times 10^{-4}$ and a Cosine Annealing schedule with 3,000 warmup steps. T4 GPU with 16GB vRAM and 29GB of system RAM was used for training in Kaggle environment with total training time of 3 hours for standard data and 12 hours for augmented data.
Table 6 summarizes the Newari speech corpus split across training, validation, and test sets, showing the number of hours and utterances for each partition.
@@
4.2.  Results and Discussion
The performance of the models was evaluated using CER as the primary metric. The experiments were conducted on the Nwāchā Munā dataset, while additional evaluations incorporated augmented data and unlabelled speech to examine the effect of data expansion and semi-supervised training. The quantitative results are summarized in Table 7.

### 30 Mar 2026 v2

@@ Abstract
Nepal Bhasha (Newari), an endangered language of the Kathmandu Valley, remains digitally marginalized due to the severe scarcity of annotated speech resources. In this work, we introduce Nwāchā Munā (/nwa:tSa: muna:/), a newly curated 5.39-hour manually transcribed Devanagari speech corpus for Nepal Bhasha, and establish the first benchmark using script-preserving acoustic modeling. We investigate whether proximal cross-lingual transfer from a geographically and linguistically adjacent language (Nepali) can rival large-scale multilingual pretraining in an ultra-low-resource Automatic Speech Recognition (ASR) setting. Fine-tuning a Nepali Conformer model reduces the Character Error Rate (CER) from a 52.54% zero-shot baseline to 17.59% with data augmentation, effectively matching the performance of the multilingual Whisper-Small model despite utilizing significantly fewer parameters. Our findings demonstrate that proximal transfer from Nepali language serves as a computationally efficient alternative to massive multilingual models. We openly release the dataset and benchmarks to digitally enable the Newari community and foster further research in Nepal Bhasha.
Keywords:  Automatic Speech Recognition, Nepal Bhasha, Speech Corpus, Low-Resource Speech Recognition
@@
Rupak Tiwari, Arju Shrestha, Rupak Raj Ghimire, Bal Krishna Bal
@@ 1.  Introduction
In the current space of AI, high-resource languages dominate in terms of resource availability while indigenous languages remain digitally marginalized. In a country like Nepal, with 124 official languages, this divide is especially vivid, creating a need for technologies like ASR to bridge the gap. Nepal Bhasha (Newari) serves as a prime example of this imbalance, despite being actively spoken by over 860,000 people in Nepal. Genetically, Nepal Bhasha belongs to the Tibeto-Burman branch of the Sino-Tibetan family Eberhard et al. (2026). Even with a six-century-old history and a designation as one of the official working languages of Bagmati Province Post Report (2024), Nepal Bhasha is currently classified by UNESCO as a ``definitely endangered'' language Moseley (2010). Building a system such as ASR empowers these communities to maintain their linguistic heritage while fully participating in the modern technological landscape.
@@
From this research, we aim to provide a scalable blueprint for other underrepresented languages in the region.
@@ 2.  Related works
One major obstacle in building ASR systems for underrepresented languages is the lack of sufficient speech data. Older methods like Hidden Markov Models required less data but struggled to capture intricate acoustic patterns. With time, new architectures like Transformers Vaswani et al. (2017) and Conformer Gulati et al. (2020) have outperformed older statistical methods, but they require larger datasets. To address this issue, researchers have increasingly turned to transfer learning. For instance, Cheng et al. (2024) demonstrated clear ASR performance gains for two at-risk Austronesian languages, Amis and Seediq, despite having minimal available speech samples. They achieved this by utilizing a novel data-selection scheme to automatically extract phonetically similar utterances from a broad multilingual corpus to augment their training data. Likewise, significant progress has been made specifically within the South Asian region. The AI4Bharat initiative Javed et al. (2022) advanced regional ASR by curating over 17,000 hours of raw speech data across 40 Indian languages for continued pre-training.
@@
In the context of Nepali ASR, early efforts relied on Hidden Markov Models (HMM) Ssarma et al. (2017) for isolated word recognition, which struggled with continuous speech. The transition to deep learning began with Regmi et al. (2019), they implemented Recurrent Neural Networks (RNN) with Connectionist Temporal Classification (CTC) loss. This was further refined by Bhatta et al. (2020) with a hybrid CNN-GRU-CTC model. Furthermore, Regmi and Bal (2021) implemented end-to-end speech recognition approach for the Nepali language. To address training instability in deeper networks, Dhakal et al. (2022) proposed a CNN-BiLSTM-ResNet architecture, achieving a CER of 17.06%. Concurrently, research began shifting towards data efficiency and cross-lingual techniques. For example, Bansal et al. (2020) explored an acoustic-phonetic approach, demonstrating that leveraging monolingual and cross-lingual information could improve performance in low-resource settings. Joshi and Shrestha (2023) applied Self-Attention Networks (SAN) to capture long-range dependencies better than recurrent models. Building on these modern architectures, Paudel et al. (2023) successfully applied a combination of CNNs and Transformers for large vocabulary continuous speech recognition.
@@
Current state-of-the-art systems have fully embraced E2E Transformer-based architectures. Poudel et al. (2025) introduced NepConformer, a Conformer-based Nepali ASR model achieving a CER of 6.01%. Recent focus has also turned to data selection, tokenization and resource efficient strategies. Ghimire et al. (2023a) proposed an Active Learning framework to select the most informative samples for speech annotation. Another work Ghimire et al. (2023b), demonstrated that using linguistically motivated sub-word units (syllables) alongside active sampling significantly outperforms standard methods. Work done by Ghimire et al. (2025) showed that using adapter-based Parameter-Efficient-Fine-Tuning (PEFT) alone improves Word Error Rate (WER) by 19% compared to Whisper-large-v2. Extending the utility of efficient adaptation, Pantha et al. (2025) demonstrated the effectiveness of PEFT for speech personalization among Nepali speakers. Most recently, Ghimire et al. (2023b) implemented new complementary loss function for addressing grammatical consistency during model training and fine-tuning. Their Rule-Based Character Constituency Loss (RBCCL) help to reduce Word Error Rates (WER) from 47.1% to 23.41% by penalizing grammatically impossible character sequences during training.
@@
While Nepali has seen consistent growth in the field of ASR, research on Nepal Bhasha remains sparse. A notable exception is the work by Meelen et al. (2024), who presented the first Newari ASR, where they achieved a CER of 12%. Their evaluation is conducted in Romanized transliteration, this fails to preserve Devanagari orthography. Their study highlighted that while transfer learning is effective, the lack of standardized orthography and limited data results in higher WER compared to Nepali. They emphasized the need for "orthography standardization" and data augmentation techniques like SpecAugment Park et al. (2019) to make ASR viable for these endangered languages.
@@
Taken together, these findings highlight a persistent gap: prior work lacks ASR systems trained on the Devanagari script using carefully curated low-resource datasets. This research clears these gaps by developing Nwāchā Munā. We utilise Devanagari script for transcribing speech of Nepal Bhasha. We also use data augmentation techniques to train robust ASR models using various cross-lingual transfer strategies and report CER for each technique.
@@ 3.1.  Data Collection Technique
The development of the Nepal Bhasha ASR required a carefully designed data collection process to ensure authenticity and reliability. Both textual and acoustic resources were gathered considering language and speaker diversity. The objective was to construct a dataset that reflects natural language usage while maintaining consistency and structural integrity. The following subsections describe the procedures adopted for textual and audio data acquisition.
@@ 3.1.1.  Textual Data Acquisition
To establish a baseline for the Nepal Bhasha ASR system, a comprehensive textual corpus was compiled from different sources of digital and print media. The data collection strategy included the use of both formal and everyday language to capture the full range of vocabulary. Digital data were extracted from Nepal Bhasha Wikipedia 1 and the OSCAR (Open Super-large Crawled Aggregated coRpus) Ortiz Suárez et al. (2020) dataset while print-based resources included regional newspapers, literary manuscripts from the Bloom Library 2, and primary school textbooks. In addition to the formal texts obtained, a targeted set of daily conversational sentences was curated manually to incorporate texts that were previously missing. The raw text was rigorously pre-processed where non-target language tokens, numerical digits, and idiosyncratic symbols were manually removed to ensure high fidelity and relevance. Table 1 shows an example of text normalization, where raw text containing Devanagari numerals and symbols is converted into a phonetic, standardized form for acoustic modeling.
Table 1: Example of Text Normalization for Nepal Bhasha ASR
Condition Sentence Structure
[…]
chagu bwasay 21% bwa du
The sentence contains 21 as a number.
[…]
chagu bwasay nīchagū sayeka bwa du
The sentence contains 21 in textual form
The overall text corpus statistics and sentence length distribution are presented in Table 2 and 3:
@@
Total Words 27,644
Unique Words in Corpus 8,599
@@ 3.1.2.  Audio Data Acquisition
Primary acoustic data was gathered through a dual-modal strategy involving original field recordings and the transcription of existing web-based audio, ensuring the model's exposure to varied acoustic environments. To maintain technical consistency across the dataset, all audio was standardized to a 16 KHz sampling rate and stored in a mono-channel WAV format, providing the necessary fidelity for deep learning-based feature extraction. For the recording phase, 18 native speakers from Banepa, Dhulikhel, Panauti, and Patan volunteered based on criteria such as native fluency and linguistic knowledge. While Newari exhibits location-based dialectal variations, the use of a standardized text across all recordings ensured that the resulting differences remained primarily tonal. To enhance the model's robustness against speaker-dependent variability, the data was stratified by age and gender. Audio recordings were captured using built-in smartphone microphones in an open environment with minimal acoustic disturbance. This field collection of 4 hours and 21 minutes of speech was further supplemented by approximately one hour of web-sourced audio (Meelen et al., 2024), which was originally in Romanized form and subsequently transliterated into Devanagari script by community members.
The overall characteristics of the Nwāchā Munā speech corpus are summarized in Tables 4-6. Table 4 presents the corpus-level statistics, including the total number of utterances, total duration, and mean utterance length. Table 5 details the age distribution of the same participant set, while Table 6 provides the gender distribution of the 18 speakers contributing to the corpus.
@@ 3.2.1.  Acoustic Modeling Strategies
To combat data scarcity in low-resource Nepal Bhasha ASR, a dual-stage data augmentation strategy was employed. By combining static (offline) and dynamic (online) methods, we introduced diverse acoustic variations to improve model robustness and prevent overfitting. This approach acts as a regularizer, forcing the model to learn invariant phonetic features rather than memorizing the specific acoustic signatures of the limited training set.
@@ 3.2.2.  Language Modeling and Decoding
In the decoder-only setup, the encoder parameters were frozen, and only the prediction network and joint network were updated during fine-tuning. By applying the phonetic similarities between Nepali and Newari, this strategy conserves the transferable acoustic representations obtained during pre-training while allowing the decoder to adapt specifically to Newari linguistic patterns.
Independent of decoder adaptation, we also integrated an external n-gram language model trained using KenLM and applied shallow fusion during inference. This introduces explicit linguistic priors to capture word-level dependencies that the limited paired data might miss. During beam search decoding, the log-probabilities from the acoustic model and the external language model were linearly combined (Toshniwal et al., 2018), with the fusion weight tuned on the development set.
@@
$y^{*}=\arg\max_{y}\left\{\log p(y\mid x)+\alpha\log p_{LM}(y)+\beta|y|_{w}\right\}$
where $\alpha$ controls the language model weight, $\beta$ is the word insertion bonus, and $|y|_{w}$ denotes the number of words in the hypothesis $y$.
Decoder-only fine-tuning improves the model’s language understanding by updating only the higher-level components while keeping the encoder fixed, which also reduces computational cost. This setup allows us to observe the effect of adapting the decoder alone in a low-resource setting. While, shallow fusion adds an external language model during decoding without changing the acoustic model itself. By evaluating these two methods independently, we can clearly measure the contributions of internal decoder adaptation and external statistical language modeling to CER, WER, and improve transcription consistency in low-resource settings.
@@ 3.2.3.  Semi-Supervised Learning
Due to the availability of limited data for Nepal Bhasha, we adopted a semi-supervised pseudo-labeling framework. An unlabeled corpus was collected from publicly available broadcast sources, including radio, podcasts, and television, and segmented into shorter utterances. To mitigate memory issues and alignment degradation in CTC-based training, excessively long utterances were discarded. The remaining audio was transcribed using our best intermediate acoustic model to generate pseudo-labels.
To ensure quality, we applied rigorous filtering. Single-word transcripts were removed, as they were typically noise-induced fragments rather than valid words. We also computed average prediction confidence scores and applied a threshold filter to reduce error reinforcement before incorporating the pseudo-labeled data into our training pipeline.
@@ 4.1.  Experimental Setup
The dataset was divided into 80% for training, 10% for validation, and 10% for testing. Due to the limited size of the dataset across 18 participants, speaker representation was maintained in all splits to prioritize overall acoustic diversity rather than enforcing strict speaker independence. Table 7 summarizes the Newari speech corpus split across training, validation, and test sets, showing the number of hours and utterances for each partition.
@@
For the Conformer model, static augmentation expanded the training set by a factor of 5 via speed perturbation ($0.9\times,1.1\times$), volume randomization, and noise injection, resulting in 23.05 total hours of audio. Dynamic augmentation was applied with a probability of $p=0.5$, involving time stretching (rate $0.8\times$$1.25\times$), pitch shifting ($\pm 4$ semitones), and Gaussian noise injection (amplitude 0.0010.0010.0150.015).
The fine-tuned Nepal Bhasha ASR model utilizes a Conformer-CTC architecture with a Byte Pair Encoding (BPE) tokenizer of vocabulary size 128. The 18-layer Conformer encoder has a d-model dimension of 256, 4 attention heads, and a convolution kernel size of 31. Regularization includes a dropout rate of 0.1 across the encoder and attention modules, alongside SpecAugment with 2 frequency masks (width 27) and 2 time masks (width 70). Optimization is performed using AdamW with a learning rate of $1\times 10^{-4}$ and a Cosine Annealing schedule with 3,000 warmup steps. T4 GPU with 16GB vRAM and 29GB of system RAM was used for training in Kaggle environment with total training time of 3 hours for standard data and 12 hours for augmented data.
For language modeling and decoding, we performed a grid search on the development set to identify optimal decoding hyperparameters across all models. We explored different configurations of language model weight ($\alpha$), word insertion bonus ($\beta$), and beam width. Based on this systematic evaluation, a 5-gram KenLM with $\alpha=0.4$, $\beta=1.5$, and a beam width of 128 was selected, as it consistently yielded the best performance improvements. The language model was trained on a text corpus consisting of approximately 51k utterances and 478k word tokens, sourced from Nepal Bhasha Wikipedia. Notably, this dataset was reserved solely for language modeling and is distinct from the audio transcription data.
For the semi-supervised learning experiments, an initial unlabeled corpus of 13.65 hours was collected. During the filtering phase, utterances longer than 15 seconds were discarded. After transcribing the audio with our intermediate acoustic model, a 70% confidence threshold was applied to the predictions to ensure pseudo-label quality. Following this rigorous filtering process, 9.33 hours of high-quality pseudo-labeled data remained for experimental training.
4.2.   Findings and Analysis
The performance of the models was evaluated using CER as the primary metric. The experiments were conducted on the Nwāchā Munā dataset, while additional evaluations incorporated augmented data and unlabelled speech to examine the effect of data expansion and semi-supervised training. The quantitative results are summarized in Table 8, where only the best-performing models at their optimal training epochs are reported.